OpenStudy (anonymous):

Please help! Will give medal! Find the first six terms of the sequence. a1 = 1, an = 4 • an-1

OpenStudy (anonymous):

a1 = 1 a2 = 4* a(2-1) a2 = 4*a1 = 4*1 = 4 a3 = 4* a(3-1) a3 = 4*a2 = 4*4 = 16 a4 = 4* a(4-1) a4 = 4*a3 = 4*16 = 64 and so on for 5th and 6th term

OpenStudy (anonymous):

thank you so much! @sangya21

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

would you be able to help me with another one if possible?

OpenStudy (anonymous):

okkk

OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -8 - 3 + 2 + 7 + ... + 67

OpenStudy (anonymous):

a = -8 d = 5 an = 67 67 = a +(n-1)d 67 = -8 +(n-1)5 75 = (n-1)*5 15 = n-1 n = 16 \[\sum_{k=0}^{n} \space= \space n/2\space (2a +(n-1)d)\] put n = 16 d = 5 a = -8

OpenStudy (anonymous):

i plugged it in but it didnt make sense

OpenStudy (anonymous):

n/2(2a +(n-1)d) 16/2 (2*-8 +(16-1)*5) 8(-16 +(15*5)) 8(-16 +75) 8*59 472